If the ace, King, queen and jack are considered to be 'face cards' in an ordinary deck, find the probability that three successive 'face cards' are drawn, without replacement, from an ordinary deck.
Answer by stanbon(75887)

If the ace, King, queen and jack are considered to be 'face cards' in an ordinary deck, find the probability that three successive 'face cards' are drawn, without replacement, from an ordinary deck.
-----------
There are 16 face card in the deck.
There are 52 cards in the deck.
P(three face) = 12C3/52C3 = (12/52)(11/51)(10/50) = 0.009954... or 11/1105
=============
Cheers,
Stan H.